ganeshie8 (ganeshie8):

prove \(43 | 6^{n+2} + 7^{2n+1} \) @ikram002p

OpenStudy (kc_kennylau):

Would it help if I convert it to \(\large 36\times6^n+7\times49^n\)?

OpenStudy (kc_kennylau):

(This kind of problem is challenging indeed I lyk dem :p)

OpenStudy (kc_kennylau):

(this is the first time i've never met this problem)

ganeshie8 (ganeshie8):

thats a good idea ! lets try to reduce further

OpenStudy (kc_kennylau):

\(\large 36\times2^n\times3^n+7\times7^n\times7^n\)? Am I going in the right way?

OpenStudy (anonymous):

@kc_kennylau that actually makes things easier. @ganeshie8 The best way to prove this is to use mathematical induction. Are you familiar with that technique?

ganeshie8 (ganeshie8):

49 --- leaves remainder 6 43

OpenStudy (kc_kennylau):

oh mathematical induction @ChristopherToni thanks

OpenStudy (kc_kennylau):

@ChristopherToni he's challenging us he actually knows how to do it

ganeshie8 (ganeshie8):

yes Chris :) nooo i dont knw the solution yet, this problem came in my friend's exam today

ganeshie8 (ganeshie8):

\(\large 36\times6^n+7\times49^n \) leaves same ramainder as \(\large 36\times6^n+7\times 6^n \)

ganeshie8 (ganeshie8):

kc_kennlau i think u nailed it. we're done but induction proof wud be beautiful to deduce if it doesnt take too many steps :)

OpenStudy (kc_kennylau):

@ganeshie8 thanks for your hint :) so it goes: \[\hspace{11pt}\large43|6^{n+2}+7^{2n+1}\]\[\large=43|36\times6^n+7\times49^n\]\[\large=43|36\times6^n+7\times6^n\]\[\large=43|43\times6^n\]\(\large=0\)? Thanks for your challenge i lyked it :)

ganeshie8 (ganeshie8):

looks neat !! thank you :)

OpenStudy (kc_kennylau):

it should be me who am thanking you :) Thanks for training my brain :D

ganeshie8 (ganeshie8):

:) induction :- n = 1 lies in the set as 43 | 6^3 + 7^3 assume n is also in the set, lets prove n+1 hmm

ganeshie8 (ganeshie8):

il wait for Chris to complete... looks it can be tricky, we need to use congruences i feel

OpenStudy (kc_kennylau):

@ganeshie8 lol sorry I thought \(\large x|y\) wasn't a statement but in fact it is so I cannot use equal sign...

ganeshie8 (ganeshie8):

to hell wid notations lol i get wat u mean exactly :p

ganeshie8 (ganeshie8):

im going for lunch... wil get back in ~1 hour and try the induction... thanks friends :)

OpenStudy (kc_kennylau):

And I'll still be here waiting :)

OpenStudy (anonymous):

By induction: The base case is \(\large n=1\). So we see that \[\large 6^{1+2} + 7^{2(1)+1} = 6^3 + 7^3 = 559 = 43\cdot 13\implies 43\mid 6^3+7^3.\]So the base case is proved. Next suppose that for some \(\large k> 1\) that \(\large 43\mid 6^{k+2} + 7^{2k+1}\) which can be written as \(\large \color{red}{6^{k+2}+7^{2k+1} = 43m}\) for some \(\large m\in\Bbb{Z}\). We seek to show that \[\large 43\mid 6^{(k+1)+2} + 7^{2(k+1)+1}=6^{k+3}+7^{2k+3}.\] So, we note that \[\large \begin{aligned} 6^{k+3} + 7^{2k+3} &= 6\cdot 6^{k+2} + 49\cdot 7^{2k+1}\\ &= 6\cdot 6^{k+2} + (43+6)\cdot 7^{2k+1}\\ &= 6\cdot 6^{k+2} + 6\cdot 7^{2k+1} + 43\cdot 7^{2k+1}\\ &= 6(\color{red}{6^{k+2}+7^{2k+1}}) + 43\cdot 7^{2k+1}\\ &= 6\cdot 43m + 43\cdot 7^{2k+1}\\ &= 43\underbrace{(6m+7^{2k+1})}_{\in\Bbb{Z}} \end{aligned} \] Hence, \(\large 6^{k+3}+7^{2k+3}\) is a multiple of 43, which now implies that \(\large 43\mid 6^{k+3}+7^{2k+3}\). Therefore, \(\large 43\mid 6^{n+2} + 7^{2n+1}\). Does this make sense, guys?

OpenStudy (anonymous):

Actually, on second thought, the base case is \(\large n=0\), since \(\large 6^2+7 = 43\) which clearly is divisible by 43.

OpenStudy (kc_kennylau):

that makes \(\Huge{\mbox{PERFECT}}\) sense :D

OpenStudy (ikram002p):

it can also solved by using concept of congruence 7^n=6mod 43 6^n=6^n mod 43 6^2=36=-7 mod 43 6^2n=-7^n mod 43 -7 = 0 mod 43 -7.7^2n=6^2(6^n) mod 43 -7^(2n+1)=6^(n+2 ) mod 43 -(7^(2n+1)+6^(n+2 ))=43k #done ^_^

OpenStudy (ikram002p):

and the funny things was it was mintion in the exam using induction is not allowed lol

ganeshie8 (ganeshie8):

Wow ! all 3 proofs look great !! xD thanks @ikram002p @kc_kennylau @ChristopherToni =))

OpenStudy (kc_kennylau):

no problem at all :)

OpenStudy (kc_kennylau):

thank you for training my brain too :)